# ComGPT: Detecting Local Community Structure with Large Language Models

### v1

@@ Introduction
dataset M R DMF_M DMF_R LCDNN LCDMD ComGPT
Dolphins 0.4685 0.4505 0.7255 0.7242 0.7558 0.7235 0.7885
Polbooks 0.5257 0.5030 0.7295 0.7286 0.7463 0.7301 0.7586
Football 0.7146 0.7107 0.8881 0.8893 0.8948 0.8134 0.8406
Amazon 0.7706 0.7554 0.8353 0.8361 0.8372 - 0.8385
DBLP 0.4185 0.4185 0.4032 0.4009 0.3152 - 0.4247
Table 1: Comparisons with other methods in terms of fscore.
@@
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms.
The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes within this network represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges.The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. In this network, nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors, and edges indicate that two authors have co-authored a paper.
Comparison algorithms
To evaluate the performance of ComGPT, we compare it with five local community detection algorithms: M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LCDMD (Guo et al. 2022). ComGPT uses gpt-3.5-turbo-0125 and sets $k_{f}$ to 0.5. Due to monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
@@
The evaluation metrics precision, recall, fscore (Powers 2020) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. fscore is the harmonic mean of precision and recall. The values of precision, recall, and fscore are between 0 and 1. The larger value implies a better algorithm performance.
@@
For the Dolphins, Polbooks, and Football datasets, each node is selected as a seed node separately. For the Amazon and DBLP datasets, We randomly select 200 nodes from the top 5,000 communities. The average fscore of all seed nodes is calculated. The detailed experimental results are shown in Table 1. Due to the unavailability of the LCDMD (Guo et al. 2022) source code, results for the Dolphins, Polbooks, and Football datasets are cited from (Guo et al. 2022), in which results for the Amazon and DBLP datasets are not available.
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LCDMD on most datasets. ComGPT is superior to other methods because GPT has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, LCDNN, and LCDMD perform better than the M method and R method. The reason why DMF_M, DMF_R, LCDNN, and LCDMD outperform the M method and R method on most datasets is that the latter two methods maximize local modularity (M and R), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, LCDNN, and LCDMD implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize dynamic membership functions to select nodes, which tend to favor nodes with large degrees. LCDNN expands the community based on the fuzzy relationships between nodes and NGC nodes. Meanwhile, LCDMD introduces a new modularity density to effectively mine the core parts of a community.
@@
Method GPT SK NSG precision recall fscore
1 NO NO NO 0.6824 0.7246 0.6978
2 YES NO YES 0.7571 0.8277 0.7832
3 YES YES NO 0.5911 0.7026 0.6221
ComGPT YES YES YES 0.8440 0.8520 0.8445
Table 2: Results of ablation experiment on the football dataset.
To explore the necessity of GPT, SK (Supplementary knowledge in graph coding), and NSG prompt in ComGPT, we conduct ablation experiments on the Football dataset with 50 randomly selected seed nodes. Three comparative models are constructed as follows: 1) Method 1 does not use GPT. The first node in $PN$ will be selected for community expansion until $PN$ is empty; 2) Method 2 only maintains GPT and NSG prompt, without SK; and 3) Method 3 only maintains GPT and SK, without NSG prompt. When calculating precision and recall, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. In this calculation, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
Table 2 shows the experimental results of Methods 1, 2, 3, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, and 3, indicating that GPT, SK, and NSG prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of precision, recall, and fscore, which demonstrates the effectiveness and importance of GPT. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3, suggesting that NSG prompt can significantly enhance its performance in local community detection.
@@
 Method Dolphins Polbooks Football
Zero-shot P 0.7195 0.5793 0.5911
R 0.3911 0.2330 0.7026
F 0.4690 0.3261 0.6221
Few-shot P 0.7136 0.5658 0.6592
R 0.3840 0.3358 0.7600
F 0.4556 0.4070 0.6894
COT P 0.7483 0.5924 0.6120
R 0.4397 0.2633 0.7205
F 0.5173 0.3572 0.6512
BAG P 0.7242 0.6005 0.6530
R 0.4704 0.2338 0.7589
F 0.5192 0.3315 0.6870
NSG P 0.9592 0.7788 0.8440
R 0.7356 0.7559 0.8520
F 0.7964 0.7506 0.8445
Table 3: Results of different prompts. P, R, and F are the abbreviations of recall, precision, and fscore, respectively.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), COT (Wei et al. 2022), and BAG (Wang et al. 2023) prompts. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the indicator values are calculated in the same way as in the ablation study.
Table 3 shows that the performance of NSG is better than that of Zero-shot, Few-shot, COT, and BAG prompts. Specifically, on Polbooks and Football datasets, the Few-shot outperforms the Zero-shot. Few-shot is worse than Zero-shot on the Dolphins dataset. This is because each dataset has a different structure. If the provided examples match the community structure characteristics of the dataset, the performance of GPT is improved. Conversely, if they do not match, GPT’s performance is diminished. The reason COT and BAG outperform Zero-shot on most datasets is that COT enhances GPT’s reasoning ability by guiding it to think step-by-step and BAG helps GPT better digest graph information. On Dolphins, Polbooks, and Football datasets, NSG outperforms other prompts. This is because the design of COT and BAG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
@@
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 13, which exhibits higher variability due to the randomness of GPT output. To confirm this, we conduct repeated experiments for $k$ = 13 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
@@ Conclusion
We explored the potential of using GPT to assist in local community detection and developed a GPT-guided methodology, termed ComGPT. Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge. Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge. Experimental results demonstrate that ComGPT outperforms the comparison algorithms. Moreover, the designed graph encoding and prompt help GPT understand the domain knowledge of graph topology and community discovery, respectively.

### v2

@@ Introduction
 Dataset M R DMF_M DMF_R LCDNN LS ComGPT
Dolphins Jaccard 0.3370 0.3128 0.6270 0.6320 0.6438* 0.7735 0.7035
Fscore 0.4673 0.4478 0.7254 0.7241 0.7558* 0.8385 0.7684
Polbooks Jaccard 0.4266 0.3986 0.6449 0.6437 0.6673 0.6919 0.6468*
Fscore 0.5184 0.4960 0.7294 0.7285 0.7463 0.7764 0.7320*
Football Jaccard 0.6117 0.6134 0.8546 0.8579 0.7547 0.2406 0.7571*
Fscore 0.6706 0.6699 0.8881 0.8893 0.8255* 0.3593 0.8103
Amazon Jaccard 0.6961 0.6771 0.7633 0.7686 0.7665* 0.6506 0.7734
Fscore 0.7706 0.7554 0.8353 0.8361* 0.8372 0.7420 0.8421
DBLP Jaccard 0.3239 0.3207* 0.2945 0.2928 0.2208 0.0650 0.3259
Fscore 0.4185 0.4185 0.4032 0.4009 0.3152 0.0850 0.4359
Table 1: Experimental results on five datasets. On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
@@
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms. The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges. The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. Nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
Baselines
To evaluate the performance of ComGPT, we compare it with M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LS (Shi et al. 2024).
Implementation Details
ComGPT uses gpt-3.5-turbo-0125. The number of potential nodes $k$ is set to 10. All experiments are conducted under Windows 10 operating system (CPU: 11th Gen Intel(R) Core(TM) i7-11700K @ 3.60GHZ, Memory: 32 GB.). Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
@@
The evaluation metrics Fscore (Powers 2011) and Jaccard (Jaccard 1912) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. Fscore is the harmonic mean of precision and recall. The values of Fscore and Jaccard are between 0 and 1. The larger value implies a better algorithm performance.
@@
For the Dolphins, Polbooks, and Football datasets, all node are selected as seed nodes. For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as seed nodes. The average Fscore and Jaccard of all seed nodes is calculated. The detailed experimental results are shown in Table 1.
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LS on most datasets. Among all comparison methods, ComGPT exhibits the best performance on the Amazon and DBLP datasets and ranks second or third on the other three datasets. Although LS achieves the best performance on the Dolphins and Polbooks datasets, its performance on the remaining three datasets is the lowest among all comparison methods. Overall, ComGPT is superior to LS. ComGPT is superior to other methods because GPT-3.5-turbo has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT-3.5-turbo to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method. The reason why DMF_M, DMF_R, and LCDNN outperform the M method and R method on most datasets is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize membership functions to select nodes with large degrees. LCDNN expands the community based on the fuzzy relations between nodes and NGC nodes.
@@
Method GPT SK NSG NSU Fscore Jaccard
1 NO NO NO NO 0.4666 0.3352
2 YES NO YES YES 0.3711 0.2414
3 YES YES NO YES 0.4855 0.3299
4 YES YES YES NO 0.7134 0.6307
ComGPT YES YES YES YES 0.7688 0.7009
Table 2: Results of ablation experiment on the Dolphins dataset.
To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins dataset with 50 randomly selected seed nodes. Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo. The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt. When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. Here, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
Table 2 shows the experimental results of Methods 1, 2, 3, 4, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the effectiveness and importance of GPT-3.5-turbo. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT-3.5-turbo to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
@@
Prompts Description
Few-shot It is given a few demonstrations of the task
at inference time as conditioning.
Zero-shot No demonstrations are allowed and only an instruction in natural language is given to the model.
CoT Let’s think step by step.
BaG Let’s construct a graph with the nodes and edges first.
Table 3: Description of the prompts.
 Prompts Dolphins Polbooks Football
Zero-shot Jaccard 0.3299 0.2191 0.5829
Fscore 0.4855 0.3490 0.6774
Few-shot Jaccard 0.3584 0.2703 0.7145
Fscore 0.5128 0.4078 0.7764
CoT Jaccard 0.3250 0.2078 0.6168
Fscore 0.4785 0.3347 0.7196
BaG Jaccard 0.2772 0.1813 0.6050
Fscore 0.4235 0.2988 0.7131
NSG Jaccard 0.7009 0.6624 0.8045
Fscore 0.7688 0.7418 0.8555
Table 4: Results of different prompts.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), CoT (Wei et al. 2022), and BaG (Wang et al. 2023) prompts, as shown in Table 3. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the metrics values are calculated in the same way as in the ablation study.
Table 4 shows that the performance of ComGPT with NSG is better than that of ComGPT with Zero-shot, Few-shot, CoT, and BaG prompts. Specifically, on Football dataset, the ComGPT with CoT and BaG outperform the ComGPT with Zero-shot. The ComGPT with CoT and BaG is worse than ComGPT with Zero-shot on Dolphins and Polbooks datasets. Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks. For instance, advanced prompts like CoT increase the number of reasoning steps for GPT-3.5-turbo, which potentially raises the likelihood of errors in the final reasoning conclusion due to mistakes in one of the intermediate steps. The reason ComGPT with Few-shot outperforms ComGPT with Zero-shot on most datasets is that we provided some examples of how the task should be executed. These examples guide the model by showing the correct process needed to solve the task. On Dolphins, Polbooks, and Football datasets, ComGPT with NSG outperforms ComGPT with other prompts. This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
@@
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output. To confirm this, we conduct repeated experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
@@ Conclusion
We explored using GPT-3.5-turbo to assist in seed expansion algorithm and developed a GPT-guided methodology, termed ComGPT. Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge. Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge. Experimental results confirm that the designed graph encoding method and prompts help GPT-3.5-turbo understand the domain knowledge and detect local community structure, respectively.